# Equivariant and Invariant Parametrized Topological Complexity

### v3

@@ 3. Equivariant parametrized topological complexity
is a $G$-fibration, see [12, Corollary 4.3].
Definition 3.1 ([12, Definition 4.1]).
@@
Suppose $\Delta\colon E\to E\times E$ is the diagonal map. Then it is clear that the image $\Delta(E)$ is a $G$-invariant subset of $E\times_{B}E$. In the next theorem, we prove the parametrized analogue of [10, Lemma 5.14] in the equivariant setting.
@@
$(1)\implies(2)$. Suppose $s\colon U\to E^{I}_{B}$ is a $G$-equivariant section of $\Pi$. Let $H\colon E^{I}_{B}\times I\to E^{I}_{B}$ be given by
@@
It is clear that $H(\gamma,t)\in E^{I}_{B}$ for all $\gamma\in E^{I}_{B}$ and $t\in I$. Hence, $H$ is well-defined. Clearly, $H$ is $G$-equivariant such that $H(\gamma,0)=\gamma$ and $H(\gamma,1)=c_{\gamma(0)}$, where $c_{e}$ is the constant path in $E$ taking the value $e\in E$. Then
$F:=\Pi\circ H\circ(s\times\mathrm{id}_{I})\colon U\times I\to E\times_{B}E$
is a $G$-homotopy such that $F_{0}=\Pi\circ\mathrm{id}_{E^{I}_{B}}\circ s=i_{U}$ and $F_{1}(u)=\Pi(H_{1}(s(u)))=\Pi(c_{s(u)(0)})=(s(u)(0),s(u)(0))\in\Delta(E)$. Hence, $F_{1}$ is the desired map.
$(2)\implies(1)$. Suppose $H\colon U\times I\to E\times_{B}E$ is a $G$-homotopy between $f$ and $i_{U}$. Let $s\colon U\to E^{I}_{B}$ be the $G$-map given by $s(u)=c_{\pi_{1}(f(u))}=c_{\pi_{2}(f(u))}$, where $\pi_{i}\colon E\times E\to E$ is the projection map onto the $i$-th factor. By $G$-homotopy lifting property of $\Pi$, there exists a $G$-homotopy $\widetilde{H}\colon U\times I\to E^{I}_{B}$ such that the following diagram
[…]
commutes. Then $\Pi\circ\widetilde{H}_{1}=H_{1}=i_{U}$ implies $\widetilde{H}_{1}$ is a $G$-equivariant section of $\Pi$ over $U$. ∎
As a consequence to the previous theorem we can now express the equivariant parametrized topological complexity as the equivariant $\Delta(E)$-LS category of the fibre product.
@@
Suppose $p\colon E\to B$ is a $G$-fibration. If $H$ and $K$ are subgroups of $G$ such that the fixed point map $p^{H}\colon E^{H}\to B^{H}$ is a $K$-fibration, then
@@
Suppose $\Pi\colon E^{I}_{B}\to E\times_{B}E$ is $G$-equivariant parametrized fibration corresponding to $p$. Then it is easily checked that
@@
by Proposition 2.7. ∎
Applying Proposition 3.4 and Corollary 2.8, we obtain the following corollary.
@@
    $\mathrm{TC}[p\colon E\to B]\leq\mathrm{TC}_{H}[p\colon E\to B]\leq\mathrm{TC}_{G}[p\colon E\to B].$
A cohomological lower bound for the equivariant parametrized topological complexity was established by the second author in [12, Theorem 4.5] using Borel cohomology. In the following theorem, we provide an alternative cohomological lower bound that is easier to compute since it relies on non-equivariant cohomology.
@@
Suppose $\mathrm{TC}_{G}[p\colon E\to B]\leq k$. Then there exists a $G$-invariant open cover $\{U_{1},\dots,U_{k}\}$ of $E\times_{B}E$ such that each $U_{i}$ admits a $G$-equivariant section of $\Pi$. By Theorem 3.2, for each $i=1,\dots,k$, there exists a $G$-homotopy $H_{i}\colon U_{i}\times I\to E\times_{B}E$ from the inclusion map $j_{U_{i}}\colon U_{i}\hookrightarrow E\times_{B}E$ to a $G$-map $f_{i}\colon U_{i}\to E\times_{B}E$ which takes values in $\Delta(E)$. Let $\overline{U_{i}}:=\rho(U_{i})$. As $I$ is locally compact, $H_{i}$ induces a homotopy $\overline{H_{i}}\colon\overline{U_{i}}\times I\to E_{B,G}$ from the inclusion map $j_{\overline{U_{i}}}\colon\overline{U_{i}}\hookrightarrow E_{B,G}$ to a map $\overline{f_{i}}\colon\overline{U_{i}}\to E_{B,G}$ which takes values in $d_{G}E$. Thus, the following diagram
[…]
is commutative. Hence, by hypothesis (1), each $u_{i}$ restricts to zero in $H^{*}(\overline{U_{i}};R)$. By long exact sequence of the pair $(E_{B,G},d_{G}E)$, there exists classes $v_{i}\in H^{*}(E_{B,G},\overline{U_{i}};R)$ such that $v_{i}$ maps to $u_{i}$ under the coboundary map $H^{*}(E_{B,G},\overline{U_{i}};R)\to H^{*}(E_{B,G};R)$. Hence, we get
@@
When $B$ is a point, we can use the above theorem to get a cohomological lower bound for $\mathrm{TC}_{G}(E)$.
The product inequality for parametrized topological complexity was proved in [8, Proposition 6.1]. We now establish the corresponding equivariant analogue.
@@
Let $\Pi_{1}\colon(E_{1})^{I}_{B_{1}}\to E_{1}\times_{B_{1}}E_{1}$ and $\Pi_{2}\colon(E_{2})^{I}_{B_{2}}\to E_{2}\times_{B_{2}}E_{2}$ be the equivariant parametrized fibrations corresponding to $p_{1}$ and $p_{2}$, respectively. If $E:=E_{1}\times E_{2}$, $B:=B_{1}\times B_{2}$ and $p:=p_{1}\times p_{2}$ is the product $(G_{1}\times G_{2})$-fibration, then it easily checked that
$E^{I}_{B}=(E_{1})^{I}_{B_{1}}\times(E_{2})^{I}_{B_{2}}\quad\text{and}\quad E\times_{B}E=(E_{1}\times_{B_{1}}E_{1})\times(E_{2}\times_{B_{2}}E_{2})$
and the $(G_{1}\times G_{2})$-equivariant parametrized fibration $\Pi\colon E^{I}_{B}\to E\times_{B}E$ corresponding to $p$ is equivalent to the product $(G_{1}\times G_{2})$-fibration
$\Pi_{1}\times\Pi_{2}\colon(E_{1})^{I}_{B_{1}}\times(E_{2})^{I}_{B_{2}}\to(E_{1}\times_{B_{1}}E_{1})\times(E_{2}\times_{B_{2}}E_{2}).$
Since a subspace of a $(G_{1}\times G_{2})$-completely normal space is itself $(G_{1}\times G_{2})$-completely normal, it follows that $(E_{1}\times_{B_{1}}E_{1})\times(E_{2}\times_{B_{2}}E_{2})$ is $(G_{1}\times G_{2})$-completely normal. Hence,
@@
Suppose $p_{i}\colon E_{i}\to B_{i}$ is a $G$-fibration for $i=1,2$. If $G$ is compact Hausdorff, then $p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration, where $G$ acts diagonally on the spaces $E_{1}\times E_{2}$ and $B_{1}\times B_{2}$. Furthermore, if $E_{1}$ and $E_{2}$ are Hausdorff, and $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
@@
In Proposition 2.14, we showed that $p_{1}\times p_{2}\colon E_{2}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration if $G$ is compact Hausdorff.
If $E_{1}$ and $E_{2}$ are Hausdorff, and $(E_{1}\times E_{1})\times(E_{2}\times E_{2})$ is completely normal, then [10, Lemma 3.12] implies that $(E_{1}\times E_{1})\times(E_{2}\times E_{2})$ is $(G\times G)$-completely normal. Hence, the desired inequality
@@
follows from Corollary 3.5 (2) and Theorem 3.7. ∎
@@
Suppose $p_{i}\colon E_{i}\to B$ is a $G$-fibration for $i=1,2$. Let $E_{1}\times_{B}E_{2}=\{(e_{1},e_{2})\in E_{1}\times E_{2}\mid p_{1}(e_{1})=p_{2}(e_{2})\}$ and let $p\colon E_{1}\times_{B}E_{2}\to B$ be the $G$-map given by $p(e_{1},e_{2})=p_{1}(e_{1})=p_{2}(e_{2})$, where $G$ acts on $E_{1}\times_{B}E_{2}$ diagonally. If $G$ is compact Hausdorff, then $p$ is a $G$-fibration. Furthermore, if $E_{1}$ and $E_{2}$ are Hausdorff, and $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
@@
where $G$ acts on $E_{i}\times E_{i}$ diagonally for $i=1,2$; and $G\times G$ acts on $(E_{1}\times E_{1})\times(E_{2}\times E_{2})$ componentwise.
@@
Note the following diagram
[…]
is a pullback in the category of $G$-spaces, where $\Delta\colon B\to B\times B$ is the diagonal map. In Corollary 2.15, we showed that $p$ is a $G$-fibration. Hence, the desired inequality
$\mathrm{TC}_{G}[p\colon E_{1}\times_{B}E_{2}\to B]$ $\leq\mathrm{TC}_{G}[p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B\times B]$
$\leq\mathrm{TC}_{G}[p_{1}\colon E_{1}\to B]+\mathrm{TC}_{G}[p_{2}\colon E_{2}\to B]-1$
follows from [12, Proposition 4.6] and Corollary 3.8. ∎
@@ 4. Invariant Parametrized Topological Complexity
That is the following diagram
@@
is a pullback. For each path $\alpha\in E^{I}_{B}$, let $b_{\alpha}$ denote the element in $B$ such that $\alpha$ take values in the fibre $p^{-1}(b_{\alpha})$. Define the map
@@
As $E^{I}_{B}\times_{E/G}E^{I}_{B}$ and $E\times_{B/G}E$ are $(G\times G)$-invariant subsets of $E^{I}_{B}\times E^{I}_{B}$ and $E\times E$ respectively, we get $(G\times G)$-action on $E^{I}_{B}\times_{E/G}E^{I}_{B}$ and $E\times_{B/G}E$, and $\Psi$ becomes a $(G\times G)$-equivariant map.
@@
$S:=\{(e_{1},e_{2},e_{3},e_{4})\in(E\times_{B}E)\times(E\times_{B}E)\mid(\gamma(1),\delta(0))\in E\times_{E/G}E\}.$
@@
We now introduce the main object of our study, which is a parametrized analogue of invariant topological complexity introduced by Lubawski and Marzantowicz in [25].
@@
The $G$-homotopy equivalence of the invariant topological complexity was established by Lubawski and Marzantowicz in [25, Proposition 2.4 and Lemma 3.8]. We will now establish the corresponding parametrized analogue. In particular, we establish the equivariant fibrewise homotopy equivalence of invariant parametrized topological complexity. We refer the reader to [12, Section 4.1] for basic information about fibrewise equivariant homotopy equivalence.
@@
Suppose $\tilde{f}=f\times f$, $\tilde{f}^{I}(\gamma,\delta)=(f\circ\gamma,f\circ\delta)$ and $\tilde{f^{\prime}}$, $\tilde{f^{\prime}}^{I}$ are defined similarly. Note that $\tilde{f}$ and $\tilde{f^{\prime}}$ are $(G\times G)$-maps. Then we have the following commutative diagram.
[…]
Since the maps $f^{\prime}\circ f$ and $\mathrm{id}_{E}$ are fibrewise $G$-homotopy equivalent, it follows that the maps $\tilde{f^{\prime}}\circ\tilde{f}$ and $\mathrm{id}_{E\times_{B/G}E}$ are $(G\times G)$-homotopy equivalent. Then, using [12, Lemma 4.10(2)], we obtain the inequality
@@
The next proposition shows that the invariant parametrized topological complexity of a $G$-fibration is a generalization of both the parametrized topological complexity of a fibration [8] and the invariant topological complexity of a $G$-space [25].
@@
(1) If $G$ acts trivially on $E$, then $\pi_{E}\colon E\to E/G$ is the identity map. Hence, $E^{I}_{B}\times_{E/G}E^{I}_{B}=E^{I}_{B}\times_{E}E^{I}_{B}$ which is homeomorphic to $E^{I}_{B}$ via the map $(\gamma,\delta)\mapsto\gamma*\delta$, where $\gamma*\delta$ is the concatenation of paths $\gamma$ and $\delta$. The inverse of this homeomorphism is given by $\alpha\mapsto\left(\left.\alpha\right|_{\left[0,\frac{1}{2}\right]},\left.\alpha\right|_{\left[\frac{1}{2},1\right]}\right)$ for $\alpha\in E^{I}_{B}$. If $G$ acts trivially on $B$, then $\pi_{B}\colon B\to B/G$ is the identity map. Hence, $E\times_{B/G}E=E\times_{B}E$. Therefore, the fibration $\Psi$ is given by
$\Psi\colon E^{I}_{B}\to E\times_{B}E,\quad\Psi(\alpha)=(\alpha(0),\alpha(1)).$
Hence, we get $\mathrm{TC}^{G}[p\colon E\rightarrow B]=\mathrm{TC}[p\colon E\rightarrow B]$.
(2) If $B=\{*\}$, then $E^{I}_{B}=E^{I}$ and $E\times_{B/G}E=E\times E$. Hence, the fibration $\Psi$ is given by
$\Psi\colon E^{I}\times_{E/G}E^{I}\to E\times E,\quad\Psi(\gamma,\delta)=(\gamma(0),\delta(1)).$
Therefore, $\mathrm{TC}^{G}[p\colon E\rightarrow\{*\}]=\mathrm{TC}^{G}(E)$. ∎
@@
Proof.
Let $E=B\times F$. Then note that $E^{I}_{B}=B\times F^{I}$ and $E\times_{B/G}E=(B\times_{B/G}B)\times(F\times F)$. As $E/G=(B\times F)/G=(B/G)\times F$, we have
where the last $G$-homeomorphism is induced by $(\gamma,\delta)\in F^{I}\times_{F}F^{I}\mapsto\gamma*\delta\in F^{I}$. Then it follows that the fibration $\Psi$ corresponding to $p$ is given by $\Psi=\mathrm{id}_{B\times_{B/G}B}\times e_{F}$, where $e_{F}\colon F^{I}\to F\times F$ is the free path space fibration corresponding to $F$. Thus, we obtain
since $G$ acts trivially on $F$. ∎
Remark 4.6.
In general, if $G$ acts non-trivially on $F$, then the equality
$\mathrm{TC}^{G}[p\colon B\times F\to B]=\mathrm{TC}^{G}(F)$
may not hold. For example, let $E=S^{1}\times S^{1}$ and $B=S^{1}$. If $G=S^{1}$ acts on $B$ by left multiplication and diagonally on $E$, then
$\mathrm{TC}^{S^{1}}[p\colon S^{1}\times S^{1}\to S^{1}]$ $=\mathrm{TC}[p/S^{1}\colon(S^{1}\times S^{1})/S^{1}\to S^{1}/S^{1}]$ by Theorem 4.29
$=\mathrm{TC}(S^{1}\to\{*\})$
$=\mathrm{TC}(S^{1})$ by Proposition 4.4
$=2.$
But $\mathrm{TC}^{S^{1}}(S^{1})=\mathrm{TC}(\{*\})=1$ by [25, Theorem 3.10].
@@
In the next theorem, we establish the parametrized analogue of [25, Lemma 3.8].
@@
$(1)\implies(2)$. Suppose $s=(s_{1},s_{2})\colon U\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ is a $(G\times G)$-equivariant section of $\Psi$. Let $H\colon(E^{I}_{B}\times_{E/G}E^{I}_{B})\times I\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ be given by

### v4

@@ 3. Equivariant parametrized topological complexity
is a $G$-fibration, see [D-EqPTC, Corollary 4.3].
Definition 3.1 ([D-EqPTC, Definition 4.1]).
@@
Suppose $\Delta\colon E\to E\times E$ is the diagonal map. Then it is clear that the image $\Delta(E)$ is a $G$-invariant subset of $E\times_{B}E$. In the next theorem, we prove the parametrized analogue of [colmangranteqtc, Lemma 5.14] in the equivariant setting.
@@
$(1)\implies(2)$. Suppose $\sigma\colon U\to E^{I}_{B}$ is a $G$-equivariant section of $\Pi$. Let $H\colon E^{I}_{B}\times I\to E^{I}_{B}$ be given by
@@
Clearly, $H$ is well-defined and $G$-equivariant, such that $H_{0}=\mathrm{id}_{E^{I}_{B}}$ and $H(\gamma,1)=c_{\gamma(0)}$, where $c_{e}$ is the constant path in $E$ taking the value $e\in E$. Then
$F:=\Pi\circ H\circ(\sigma\times\mathrm{id}_{I})\colon U\times I\to E\times_{B}E$
is a $G$-homotopy such that $F_{0}=i_{U}$ and $F_{1}(u)=(\sigma(u)(0),\sigma(u)(0))\in\Delta(E)$. Hence, $F_{1}$ is the desired map.
$(2)\implies(1)$. Suppose $H\colon U\times I\to E\times_{B}E$ is a $G$-homotopy between $f$ and $i_{U}$. Let $\sigma\colon U\to E^{I}_{B}$ be the $G$-map given by $\sigma(u)=c_{\pi_{1}(f(u))}=c_{\pi_{2}(f(u))}$, where $\pi_{i}\colon E\times_{B}E\to E$ is the projection map onto the $i$-th factor. Since $\Pi$ is a $G$-fibration, there exists a $G$-homotopy $\widetilde{H}\colon U\times I\to E^{I}_{B}$ such that $\widetilde{H}_{0}=\sigma$ and $\Pi\circ\widetilde{H}=H$. Then $\Pi\circ\widetilde{H}_{1}=H_{1}=i_{U}$ implies $\widetilde{H}_{1}$ is a $G$-equivariant section of $\Pi$ over $U$. ∎
As a consequence of the previous theorem, we can now express the equivariant parametrized topological complexity as the equivariant $\Delta(E)$-LS category of the fibre product.
@@
Suppose $p\colon E\to B$ is a $G$-fibration. If $H$ and $K$ are subgroups of $G$ such that $E^{H}$ and $B^{H}$ are $K$-invariant, and the fixed point map $p^{H}\colon E^{H}\to B^{H}$ is a $K$-fibration, then
@@
Suppose $\Pi\colon E^{I}_{B}\to E\times_{B}E$ is the $G$-equivariant parametrized fibration corresponding to $p$. Then it is easily checked that
@@
by Proposition˜2.12. ∎
Applying Proposition˜3.4 and Proposition˜2.12, we obtain the following corollary.
@@
    $\mathrm{TC}[p\colon E\to B]\leq\mathrm{TC}_{K}[p\colon E\to B]\leq\mathrm{TC}_{G}[p\colon E\to B].$
A cohomological lower bound for the equivariant parametrized topological complexity was established by the second author in [D-EqPTC, Theorem 4.5] using Borel cohomology. In the following theorem, we provide an alternative cohomological lower bound based on ordinary cohomology, which should, in principle, be easier to compute. The proof follows arguments similar to those in [grant2019symmetrized, Theorem 4.6].
@@
Suppose $\mathrm{TC}_{G}[p\colon E\to B]\leq k$. Then there exists a $G$-invariant open cover $\{U_{1},\dots,U_{k}\}$ of $E\times_{B}E$ such that each $U_{i}$ admits a $G$-equivariant section of $\Pi$. By Theorem˜3.2, for each $i=1,\dots,k$, there exists a $G$-homotopy $H_{i}\colon U_{i}\times I\to E\times_{B}E$ from the inclusion map $j_{U_{i}}\colon U_{i}\hookrightarrow E\times_{B}E$ to a $G$-map $f_{i}\colon U_{i}\to E\times_{B}E$ which takes values in $\Delta(E)$. Let $\overline{U_{i}}:=\rho(U_{i})$. As $I$ is locally compact, $H_{i}$ induces a homotopy $\overline{H_{i}}\colon\overline{U_{i}}\times I\to E_{B,G}$ from the inclusion map $j_{\overline{U_{i}}}\colon\overline{U_{i}}\hookrightarrow E_{B,G}$ to a map $\overline{f_{i}}\colon\overline{U_{i}}\to E_{B,G}$ which takes values in $d_{G}E$. Thus, the following diagram
$d_{G}E$$\overline{U_{i}}$$E_{B,G}$$\scriptstyle{j_{dX}}$$\scriptstyle{j_{\overline{U_{i}}}}$$\scriptstyle{\overline{f_{i}}}$
is commutative. Hence, by hypothesis (1), each $u_{i}$ restricts to zero in $H^{*}(\overline{U_{i}};R)$. By the long exact sequence of the pair $(E_{B,G},d_{G}E)$, there exist classes $v_{i}\in H^{*}(E_{B,G},\overline{U_{i}};R)$ such that $v_{i}$ maps to $u_{i}$ under the coboundary map $H^{*}(E_{B,G},\overline{U_{i}};R)\to H^{*}(E_{B,G};R)$. Hence, we get
@@
When $B$ is a point, the above theorem yields a cohomological lower bound for $\mathrm{TC}_{G}(E)$, as illustrated in the example below.
Example 3.7.
Let $G=\mathbb{Z}_{2}$ act on $S^{n}$ via the antipodal action. Then $(E\times E)/\mathbb{Z}_{2}$ is the projective product space in the sense of Davis [Davis]. If $n$ is odd, the mod-2 cohomology ring of $(E\times E)/\mathbb{Z}_{2}$ is given by
$\frac{\mathbb{Z}_{2}[y]}{\langle y^{n+1}\rangle}\otimes\Lambda_{\mathbb{Z}_{2}}[x],\quad\text{with }|y|=1,\,|x|=n\text{ (cohomological degrees)},$
where $\mathbb{Z}_{2}[y]/\langle y^{n+1}\rangle$ is the mod-2 cohomology ring of $d_{\mathbb{Z}_{2}}(S^{n})=\mathbb{R}P^{n}$. If $n$ is even, the mod-2 cohomology ring of $(E\times_{B}E)/\mathbb{Z}_{2}$ is the same as above with the additional relation $x^{2}=y^{n}x$, see [Davis, Theorem 2.1]. Hence, the cohomology classes of $(E\times E)/\mathbb{Z}_{2}$ that restrict to zero on $d_{Z_{2}}(S^{n})$ are precisely those in $\Lambda_{\mathbb{Z}_{2}}[x]$. We note that:
  • If $n$ is odd, then $x^{2}=0$; see Equation (2.5) in [Davis, Theorem 2.1].
  • If $n$ is even, then $x^{3}=(y^{n}x)x=y^{n}x^{2}=y^{n}(y^{n}x)=y^{2n}x=0$, since $y^{n+1}=0$.
Hence, by Theorem˜3.6, the lower bound on $\mathrm{TC}_{\mathbb{Z}_{2}}(S^{n})$ is 2 if $n$ is odd and 3 if $n$ is even. We note that
$\mathrm{TC}_{\mathbb{Z}_{2}}(S^{n})=\begin{cases}2,&\text{for }n\text{ odd},\\ 3,&\text{for }n\text{ even},\end{cases}$
see [eqtcprodineq, Corollary 4.4]. Hence, the lower bound we obtain is optimal. We note that we can obtain the same lower bounds using the inequality $\mathrm{TC}(S^{n})\leq\mathrm{TC}_{\mathbb{Z}_{2}}(S^{n})$.
The product inequality for parametrized topological complexity was proved in [farber-para-tc, Proposition 6.1]. We now establish the corresponding equivariant analogue.
@@
Let $\Pi_{1}$ and $\Pi_{2}$ denote the equivariant parametrized fibrations corresponding to $p_{1}$ and $p_{2}$, respectively. Then the equivariant parametrized fibration corresponding to the $(G_{1}\times G_{2})$-fibration $p_{1}\times p_{2}$ is equivalent to the product $(G_{1}\times G_{2})$-fibration $\Pi_{1}\times\Pi_{2}$, see [farber-para-tc, Proposition 6.1] for this identification. Since a subspace of a $(G_{1}\times G_{2})$-completely normal space is itself $(G_{1}\times G_{2})$-completely normal, it follows that $(E_{1}\times_{B_{1}}E_{1})\times(E_{2}\times_{B_{2}}E_{2})$ is $(G_{1}\times G_{2})$-completely normal. Hence,
@@
Suppose $p_{i}\colon E_{i}\to B_{i}$ is a $G$-fibration for $i=1,2$. If $G$ is compact Hausdorff, then $p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration, where $G$ acts diagonally on the spaces $E_{1}\times E_{2}$ and $B_{1}\times B_{2}$. Furthermore, if $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
@@
In Proposition˜2.18, we showed that $p_{1}\times p_{2}$ is a $G$-fibration. Moreover, by [colmangranteqtc, Lemma 3.12], it follows that $(E_{1}\times E_{1})\times(E_{2}\times E_{2})$ is $(G\times G)$-completely normal. Hence, the desired inequality
@@
follows from Corollary˜3.5 (2) and Theorem˜3.8. ∎
@@
Suppose $p_{i}\colon E_{i}\to B$ is a $G$-fibration for $i=1,2$. Let $E_{1}\times_{B}E_{2}=\{(e_{1},e_{2})\in E_{1}\times E_{2}\mid p_{1}(e_{1})=p_{2}(e_{2})\}$ and let $p\colon E_{1}\times_{B}E_{2}\to B$ be the $G$-map given by $p(e_{1},e_{2})=p_{1}(e_{1})=p_{2}(e_{2})$, where $G$ acts on $E_{1}\times_{B}E_{2}$ diagonally. If $G$ is compact Hausdorff, then $p$ is a $G$-fibration. Furthermore, if $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
@@
In Corollary˜2.19, we established that $p$ is a $G$-fibration. Then, by identifying $B$ with its image under the diagonal map $\Delta\colon B\to B\times B$ in the base of the fibration $p_{1}\times p_{2}$ (see (5)), the desired inequality follows from [D-EqPTC, Proposition 4.6] and Corollary˜3.9. ∎
@@ 4. Invariant Parametrized Topological Complexity
Then the following diagram
@@
is a pullback. For each path $\alpha\in E^{I}_{B}$, let $b_{\alpha}$ denote the element in $B$ such that $\alpha$ takes values in the fibre $p^{-1}(b_{\alpha})$. Define the map
@@
As $E^{I}_{B}\times_{E/G}E^{I}_{B}$ and $E\times_{B/G}E$ are $(G\times G)$-invariant subsets of $E^{I}_{B}\times E^{I}_{B}$ and $E\times E$ respectively, we get a $(G\times G)$-action on $E^{I}_{B}\times_{E/G}E^{I}_{B}$ and $E\times_{B/G}E$, and $\Psi$ becomes a $(G\times G)$-equivariant map.
@@
$S:=\{(e_{1},e_{2},e_{3},e_{4})\in(E\times_{B}E)\times(E\times_{B}E)\mid(e_{2},e_{3})\in E\times_{E/G}E\}.$
@@
We now introduce the main object of our study, which is a parametrized analogue of invariant topological complexity introduced by Lubawski and Marzantowicz in [invarianttc].
@@
The $G$-homotopy invariance of the invariant topological complexity was established by Lubawski and Marzantowicz in [invarianttc, Proposition 2.4 and Lemma 3.8]. We will now establish the corresponding parametrized analogue. In particular, we establish the equivariant fibrewise homotopy invariance of invariant parametrized topological complexity. We refer the reader to [D-EqPTC, Section 4.1] for basic information about fibrewise equivariant homotopy equivalence.
@@
Suppose $\tilde{f}=f\times f$, $\tilde{f}^{I}(\gamma,\delta)=(f\circ\gamma,f\circ\delta)$ and $\tilde{f^{\prime}}$, $(\tilde{f^{\prime}})^{I}$ are defined similarly. Note that $\tilde{f}$ and $\tilde{f^{\prime}}$ are $(G\times G)$-maps. Then we have the following commutative diagram.
$E^{I}_{B}\times_{E/G}E^{I}_{B}$$(E^{\prime})^{I}_{B}\times_{E^{\prime}/G}(E^{\prime})^{I}_{B}$$E^{I}_{B}\times_{E/G}E^{I}_{B}$$E\times_{B/G}E$$E^{\prime}\times_{B/G}E^{\prime}$$E\times_{B/G}E.$$\scriptstyle{\tilde{f}^{I}}$$\scriptstyle{\Psi}$$\scriptstyle{\Psi^{\prime}}$$\scriptstyle{(\tilde{f^{\prime}})^{I}}$$\scriptstyle{\Psi}$$\scriptstyle{\tilde{f}}$$\scriptstyle{\tilde{f^{\prime}}}$
Since the maps $f^{\prime}\circ f$ and $\mathrm{id}_{E}$ are fibrewise $G$-homotopy equivalent, it follows that the maps $\tilde{f^{\prime}}\circ\tilde{f}$ and $\mathrm{id}_{E\times_{B/G}E}$ are $(G\times G)$-homotopy equivalent. Then, using [D-EqPTC, Lemma 4.10 (2)], we obtain the inequality
@@
The next proposition shows that the invariant parametrized topological complexity of a $G$-fibration is a generalization of both the parametrized topological complexity of a fibration [farber-para-tc] and the invariant topological complexity of a $G$-space [invarianttc].
@@
  3. (3)
    If $E=B\times F$ and $p\colon B\times F\to B$ is the trivial $G$-fibration with $G$ acting trivially on $F$, then $\mathrm{TC}^{G}[p\colon B\times F\to B]=\mathrm{TC}(F).$
@@
(1) If $G$ acts trivially on $E$, then $E^{I}_{B}\times_{E/G}E^{I}_{B}=E^{I}_{B}\times_{E}E^{I}_{B}$. We note that the map
$E^{I}_{B}\to E^{I}_{B}\times_{E}E^{I}_{B},\quad\alpha\mapsto\left(\left.\alpha\right|_{\left[0,\frac{1}{2}\right]},\left.\alpha\right|_{\left[\frac{1}{2},1\right]}\right),$
is a homeomorphism, whose inverse is given by concatenation of paths. If $G$ acts trivially on $B$, then $E\times_{B/G}E=E\times_{B}E$. Therefore, the corresponding fibration $\Psi$ is given by (8). Hence, we get $\mathrm{TC}^{G}[p\colon E\rightarrow B]=\mathrm{TC}[p\colon E\rightarrow B]$.
(2) If $B=\{*\}$, then $E^{I}_{B}=E^{I}$ and $E\times_{B/G}E=E\times E$. Hence, the corresponding fibration $\Psi$ is given by (7). Therefore, $\mathrm{TC}^{G}[p\colon E\rightarrow\{*\}]=\mathrm{TC}^{G}(E)$.
(3) We note that $E^{I}_{B}=B\times F^{I}$ and $E\times_{B/G}E=(B\times_{B/G}B)\times(F\times F)$. Since $E/G=(B\times F)/G=(B/G)\times F$, it follows that
where the last $G$-homeomorphism is induced by $(\gamma,\delta)\in F^{I}\times_{F}F^{I}\mapsto\gamma*\delta\in F^{I}$. Consequently, the corresponding fibration $\Psi$ is given by $\Psi=\mathrm{id}_{B\times_{B/G}B}\times e_{F}$, where $e_{F}\colon F^{I}\to F\times F$ denotes the free path space fibration of $F$. Hence,
since $G$ acts trivially on $F$. ∎
@@
In the next theorem, we establish the parametrized analogue of [invarianttc, Lemma 3.8].
@@
$(1)\implies(2)$. Suppose $\sigma=(\sigma_{1},\sigma_{2})\colon U\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ is a $(G\times G)$-equivariant section of $\Psi$. Let $H\colon(E^{I}_{B}\times_{E/G}E^{I}_{B})\times I\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ be given by